OpenStudy (anonymous):

@phi @Luigi0210 @Compassionate @cwrw238 @amistre64 @thomaster can you help me with something?

OpenStudy (anonymous):

Describe a method you can use to find the area of the following shape. Provide specific details in your answer, including coordinates and descriptions of shapes used. Then, find the actual area of the figure.

OpenStudy (phi):

You can always divide a shape into triangles and find the area of each triangle. However, the math gets a bit messy. In this problem we can form a rectangle around the shape. It's easy to find the area of the rectangle: width * height then notice we have 4 triangles that are "cut out" of the rectangle. all of them are easy to figure out. 1/ 2 base *height

OpenStudy (anonymous):

I'm having a hard time visualizing that.

OpenStudy (anonymous):

making triangles will be your best option, and maybe a large square in the middle. so use the triangle, and square area formulas.

OpenStudy (anonymous):

Can someone draw it for me? I'm still confused.

OpenStudy (phi):

find the area of the "box" as the first step

OpenStudy (phi):

The box is a square, with width 8 and height 8 (you can count these out) what is its area ?

OpenStudy (anonymous):

Ok one sec let me find the area

OpenStudy (anonymous):

The area is 64

OpenStudy (phi):

yes. now find the area of the top left triangle that is cut out of the square count out the length of its base and height. then use 1/ 2 * base * height do that for top right triangle and the bottom right triangle for the bottom left triangle, its base is AE and its height is from the line AE to the point F

OpenStudy (anonymous):

Top left is 7.5. working on the rest

OpenStudy (anonymous):

http://www.sketchtoy.com/63125350

OpenStudy (anonymous):

@robtobey @IMStuck can you help me with the rest?

OpenStudy (imstuck):

what is 'the rest'? What do you not have an answer for?

OpenStudy (anonymous):

He told me to find the area of the box, and then the area for the triangles in the box, but I don't know what to do next.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@tkhunny @dan815 @nincompoop

OpenStudy (tkhunny):

You've an 8x8 square. The area is 64 units^2 What is the area of the four white triangles?

OpenStudy (anonymous):

The areas are 7.5, 5,5, and 3 of the four white triangles.

OpenStudy (tkhunny):

Why are there only 3 numbers?

OpenStudy (anonymous):

there are four. 1 is 7.5, 2 is 5, 3 is 5, 4 is 3

OpenStudy (anonymous):

two of them have an area of 5, unless i got it incorrect. :$

OpenStudy (tkhunny):

Okay, now subtract those from 64 and you should be done.

OpenStudy (anonymous):

Okay, one second.

OpenStudy (tkhunny):

BTW I'm not seeing a 7.5. That looks like 9 to me. Try again.

OpenStudy (anonymous):

I should use the areas as 9, 5, 5, and 3?

OpenStudy (anonymous):

Is the answer 42?

OpenStudy (tkhunny):

Why are you asking me? Did you do it correctly? Did you make a deliberate error? No guessing!! Perform whatever is needed with confidence. Ask yourself. Is it right?

OpenStudy (anonymous):

I did it the way you told me to, and I think it is correct. I was just curious if it was indeed correct.

OpenStudy (anonymous):

I don't guess.

OpenStudy (phi):

Yes, I get square is 64 the top left triangle is 1/ 2 * 3*6 = 9 top right is 1/ 2 * 2*5 = 5 bottom right is 1/ 2 * 2 *3= 3 bottom left is 1/ 2 * 5*2 = 5 sum of "cut outs" is 9+5+3+5= 22 area of "shape" is 64 - 22 = 42 which matches what you got.

OpenStudy (anonymous):

Yay! Thank you. @phi to the rescue, once again!

OpenStudy (tkhunny):

No, phi did not rescue you. You rescued yourself and didn't find the courage to see it. THAT will get you there, not phi or tkhunny. YOU confirm it.

OpenStudy (anonymous):

Thank you @tkhunny